# spade: Synthesizing Assertions for Large Language Model Pipelines

### v1

@@ Abstract.
Operationalizing large language models (LLMs) for custom, repetitive data pipelines is challenging, particularly due to their unpredictable and potentially catastrophic failures. Acknowledging the inevitability of these errors, we focus on identifying when LLMs may be generating incorrect responses when used repeatedly as part of data generation pipelines. We present spade, a method for automatically synthesizing assertions that identify bad LLM outputs. spade analyzes prompt version histories to create candidate assertion functions and then selects a minimal set that fulfills both coverage and accuracy requirements. In testing across nine different real-world LLM pipelines, spade efficiently reduces the number of assertions by 14% and decreases false failures by 21% when compared to simpler baselines.
@@ 1. Introduction
Developer Version History $\mathcal{P}_{\_}0,\ldots,\mathcal{P}_{\_}i$ All Examples Current Prompt Template $\mathcal{P}_{\_}{i+1}$ LLM Response $\spadesuit$ spadepteditsptformatsand sendsptsaves toptlabelsptgeneratesptsaves toptanalyzes $\Delta\mathcal{P}$s to generatecandidate assertionsptfilters assertionsbased on coverageand accuracy
@@
Figure 1. Before a developer deploys a prompt template to production, spade analyzes the deltas (i.e., diffs) between consecutive prompt templates to generate assertions and uses a handful of labeled responses to filter out redundant and inaccurate assertions while maintaining coverage.
There is a lot of excitement around the use of large language models (LLMs) in data pipelines (Fernandez et al., 2023). Much of the enthusiasm is due to their simplicity: without needing large labeled datasets, one can easily create an automated, intelligent pipeline that executes any arbitrary data generation task within seconds—simply by prompting an LLM with natural language. However, operationalizing LLMs for unsupervised, repeated, or large-scale data generation tasks presents significant challenges (Parameswaran et al., 2023), such as potential errors, inappropriate responses, or hallucinations (Zamfirescu-Pereira et al., 2023; Shankar et al., 2022a).
@@
Consider a movie streaming platform using LLMs to generate personalized movie recommendation notes. A developer might write a prompt template like: “write a personalized note for why a user should watch {movie_name} given the following information about the user: {personal_info} to be executed for multiple user-movie pairs. The templatized variables represent information that would get injected into the pipeline at runtime, allowing the pipeline to generate data for a variety of inputs. In theory, this prompt seems adequate, but the developer might observe some issues while testing it on a few example inputs: the LLM might reference a movie the user never watched, cite a sensitive attribute (e.g., race or ethnicity), or even have a basic issue like an overly short response.
To deal with such shortcomings, frameworks incorporating assertions into LLM pipelines have been proposed to filter out bad generations before reaching end-users (Rebedea et al., 2023; Singhvi et al., 2023). Given the unavoidable nature of LLM errors (Kalai and Vempala, 2023), such assertions are key to successful deployments of LLM pipelines that generate data. However, developers find it tough to write assertions for custom LLM pipelines (Parnin et al., 2023). Challenges include predicting all possible LLM failure modes, the time-consuming nature of writing assertions with various specification methods (like Python functions or LLM calls), the necessity for precision in assertions (especially those involving LLM calls), and the fact that many LLM pipeline developers lack software engineering expertise or coding experience (Kim et al., 2023; Zamfirescu-Pereira et al., 2023). Moreover, if there are too many or non-informative assertions, developers can be overwhelmed monitoring their results.
In this paper, we identify a new problem of automatically generating a good set of assertions for any data generation pipeline that leverages LLMs. While an assertion returns true (i.e., success) or false (i.e., failure) for an LLM response, a set of assertions returns the conjunction of individual assertions. A “good” set of assertions has minimal overlap and allows the developer to trade off between false successes and false failures. We decompose the problem into two components—candidate assertion generation and filtering—and present spade (System for Prompt Analysis and Delta-Based Evaluation, Figure 1). For candidate assertion generation, one may consider directly querying an LLM to “write assertions for x prompt,” but this may not cover the criteria the developer wants. We propose generating candidates from prompt deltas (i.e., diffs between two consecutive prompt versions), which often indicate specific failure modes of LLMs. For example, a developer might add an instruction like “avoid flowery language,” motivating an assertion to check the response language. We present an analysis of 19 custom data generation pipelines from users from LangChain (a startup that helps people build LLM pipelines) and pipeline prompt version histories, forming a taxonomy of prompt deltas. spade first automatically categorizes prompt deltas within the taxonomy, then synthesizes Python functions (that may include LLM calls) as candidate assertions. We demonstrate the potential of these assertions with a public release of this component of spade, with over 1300 uses across more than 10 sectors like finance, medicine, and IT (Shankar et al., 2023b).
Our analysis of candidate assertions revealed redundancy, inaccuracy, and a large number of functions, often exceeding 50 for just a few prompt deltas. Redundancy stems from repeated refinements to similar portions of a prompt or prompts with ambiguous instructions (e.g., “return a concise response”). Reducing this redundancy is not straightforward, even for engineers, since assertions can involve LLM calls with varying accuracies, motivating an automated filtering component. One approach is to use developer-labeled LLM responses to estimate each assertion’s false failure rate (FFR) and eliminate each one exceeding a developer-defined FFR threshold. However, the remaining assertions can have an FFR that cumulatively exceeds this threshold, and redundancies may persist. We show that selecting a small subset of assertions to meet failure coverage and FFR criteria is NP-hard. That said, we may express the problem as an integer linear program (ILP) and use an ILP solver to identify a solution. However, we find that labeled responses may not represent all failure modes, leading to omission of valuable assertions. For instance, in our movie recommendation scenario, an assertion that correctly verifies if the LLM-generated note is under 200 words will get discarded if all responses in our developer-labeled sample are below this limit. To expand coverage, active learning and weak supervision approaches can be used to sample and label new LLM prompt-response pairs for each candidate assertion (Castro and Nowak, 2008; Ratner et al., 2020), but this may be expensive or inaccessible for non-programmers. We introduce assertion subsumption as a way of ensuring comprehensive coverage: if one assertion doesn’t encompass the failure modes of another, both are selected. As such, spade selects a minimal set of assertions that respects failure coverage, accuracy, and subsumption constraints. Overall, we make the following contributions:
Prompt Deltas for Candidate Assertion Generation (Section 2). We identify prompt version history as a rich source of correctness criteria for data-generating LLM pipelines. We infer a taxonomy of assertion criteria based on prompt deltas from 19 LLM pipelines with version history, which we publish. We publicly released a tool [^1] https://spade-beta.streamlit.app to generate candidate assertions for any pipeline, providing early evidence for the utility of auto-generated assertions with over 1300 deployments while identifying opportunities to filter these assertions.
Filtering Method that Requires Little Data (Section 3). We show that selecting a minimal set of assertions while covering failure modes and meeting accuracy requirements is NP-hard. We express this problem as an ILP. Given that assertions may be redundant and inaccurate, and that LLM developers may not have engineering expertise or enough data to use existing data-driven approaches to ML pipeline testing (e.g., training models), we introduce assertion subsumption as a novel proxy for coverage in low-data settings.
@@
Empirical Study (Section 4). We present spade, our system that automatically generates assertions for data-generating LLM pipelines. For nine LLM pipelines with prompt version history, we collect labeled prompt-response pairs (eight of which we open-source). spade  [^2] https://github.com/shreyashankar/spade-experiments generates assertions with good failure coverage and few false failures (i.e., failing a good response) across all pipelines. Our subsumption-based ILP outperforms simpler baselines that do not consider interactions between assertions by reducing assertions by 14% and lowering the false failure rate by 21%.
@@ 2. Identifying Candidate Assertions
Version $i$
$\Delta\mathcal{P}_{\_}i$
Possible New Assertion Criteria
1
+ Write a personalized note for why a user should watch {movie_name} given the following information about the user: {personal_info}.
Response should be personalized and relevant to the given user information
2
+ Include elements from the movie’s genre, cast, and themes that align with the user’s interests.
Response includes specific references to the user’s interests related to the movie’s genre, cast, and themes
3
+ Ensure the recommendation note is concise.
Response should be concise
4
- Ensure the recommendation note is concise. + Ensure the recommendation note is concise, not exceeding 100 words.
Response should be within the 100 word limit
5
- Include elements from the movie’s genre, cast, and themes that align with the user’s interests. + Mention the movie’s genre and any shared cast members between the {movie_name} and other movies the user has watched.
Response should mention genre and verify cast members are accurate
6
+ Mention any awards or critical acclaim received by movie_name.
Response should include references to awards or critical acclaim of the movie
7
+ Do not mention anything related to the user’s race, ethnicity, or any other sensitive attributes.
Response should not include references to sensitive personal attributes
Table 1. Comparison of 7 prompt versions for an LLM pipeline to write personalized movie recommendations. In each $\Delta\mathcal{P}_{\_}i$, a sentence starts with “+” if it is a newly added sentence; a sentence starts with “-” if it is removed from $\mathcal{P}_{\_}{i-1}$.
@@ 2.1. Prompt Deltas
A single-step LLM pipeline consists of a prompt template $\mathcal{P}$, which is formatted with an input tuple $t$ to derive a prompt $p$, and returns a response $r$. There can be many versions of $\mathcal{P}$, depending on how a developer iterates on their prompt template. Let $\mathcal{P}_{\_}0$ be the empty string, the 0th version, and let $\mathcal{P}_{\_}i$ be the ith version of a template. In Section 1’s movie recommendation examples, suppose there are 7 versions, where $\mathcal{P}_{\_}7$ is the following: “Write a personalized note for why a user should watch {movie_name} given the following information about the user: {personal_info}. Ensure the recommendation note is concise, not exceeding 100 words. Mention the movie’s genre and any shared cast members between the {movie_name} and other movies the user has watched. Mention any awards or critical acclaim received by {movie_name}. Do not mention anything related to the user’s race, ethnicity, or any other sensitive attributes.”
We define a prompt delta $\Delta\mathcal{P}_{\_}{i+1}$ to be the diff between $\mathcal{P}_{\_}i$ and $\mathcal{P}_{\_}{i+1}$. Concretely, a prompt delta $\Delta\mathcal{P}$ is a set of sentences, where each sentence is tagged as an addition (i.e., “+”) or deletion (i.e., “-”). For example, Table 1 shows the $\Delta\mathcal{P}$s for a number of versions. Each sentence in $\Delta\mathcal{P}_{\_}i$ is composed of additions (i.e., new sentences in $\mathcal{P}_{\_}i$ that didn’t exist in $\mathcal{P}_{\_}{i-1}$) and deletions (i.e., sentences in $\mathcal{P}_{\_}{i-1}$ that don’t exist in $\mathcal{P}_{\_}{i}$). A modification to a sentence is represented by a deletion and addition—for example, $\Delta\mathcal{P}_{\_}{6}$ in Table 1 contains some new instructions added to a sentence from $\mathcal{P}_{\_}{5}$. Each addition in $\Delta\mathcal{P}_{\_}{i}$ indicates possible assertion criteria, as shown in the right-most column of Table 1.
@@ 2.2. Prompt Delta Analysis
We analyzed 19 LLM pipelines collected from LangChain users, each of which consists of between 3 and 11 historical prompt template versions. These pipelines span various tasks, from generating workout summaries to custom question-and-answering chatbots. Table 6 in Appendix B shows a summary of the pipelines, including a description of each pipeline and the number of prompt versions. For each pipeline, we categorized prompt deltas, i.e., $\Delta\mathcal{P}_{\_}i$, into different types—for example, instructing the LLM to include a new phrase in each response (i.e., inclusion), or instructing the LLM to respond with a certain tone (i.e., qualitative criteria). We iterated on our categories 4 times, ultimately producing the taxonomy in Figure 2. The taxonomy-annotated dataset of prompt versions can be found online [^3] https://github.com/shreyashankar/spade-experiments/blob/main/taxonomylabels.csv. Most prompt deltas fell under Data Integration (i.e., adding placeholder variables) and Workflow Description (i.e., adding instructions for the LLM to perform the task more accurately). In Table 2, we show sample prompt deltas for each category in our taxonomy, using the movie recommendation pipeline example from the introduction.
@@
Response Format Instruction
Structure guidelines.
“+ Start response with ‘You might like…’.”
Example Demonstration
Illustrative example.
“+ For example, here is a response for sci-fi fans…”
Prompt Clarification
Refines prompt/removes ambiguity.
“- Discuss/+ Explain movie fit…”
Workflow Description
Describe “thinking” process.
“+ First, analyze viewing history…”
Data Integration
Adds placeholders.
“+ Include user’s {genre} reviews.”
Quantity Instruction
Adds numerical content.
“+ Keep note under 100 words.”
Inclusion Instruction
Directs specific content.
“+ Mention movie awards/acclaim.”
Exclusion Instruction
Advises on omissions.
“+ Avoid movie plot spoilers.”
Qualitative Criteria
Sets stylistic attributes.
“+ Maintain friendly, positive tone.”
@@
When deriving assertion candidates from automatically-identified prompt deltas, assertion quality clearly depends on the accuracy of the identified categories. Therefore, we confirmed GPT-4’s correct categorization of prompt deltas (as of October 2023). We assigned ground truth categories to all prompt versions from the 19 pipelines, and GPT-4 achieved an F1 score of 0.8135. The prompt used for category extraction from prompt deltas is detailed in Appendix C.

### v2

# spade: Synthesizing Data Quality Assertions for Large Language Model Pipelines

@@ Abstract.
Large language models (LLMs) are being increasingly deployed as part of pipelines that repeatedly process or generate data of some sort. However, a common barrier to deployment are the frequent and often unpredictable errors that plague LLMs. Acknowledging the inevitability of these errors, we propose data quality assertions to identify when LLMs may be making mistakes. We present spade, a method for automatically synthesizing data quality assertions that identify bad LLM outputs. We make the observation that developers often identify data quality issues during prototyping prior to deployment, and attempt to address them by adding instructions to the LLM prompt over time. spade therefore analyzes histories of prompt versions over time to create candidate assertion functions and then selects a minimal set that fulfills both coverage and accuracy requirements. In testing across nine different real-world LLM pipelines, spade efficiently reduces the number of assertions by 14% and decreases false failures by 21% when compared to simpler baselines. spade has been deployed as an offering within LangSmith, LangChain’s LLM pipeline hub, and has been used to generate data quality assertions for over 2000 pipelines across a spectrum of industries.
@@ 1. Introduction
$\spadesuit$ spade Developer Current Prompt Template $\mathcal{P}_{\_}{i+1}$ Version History $\mathcal{P}_{\_}0,\ldots,\mathcal{P}_{\_}i$ LLM Output All Example Inputs and Outputs 1 Generate Assertion Candidates 2 Filter Bad Assertions Assertions to Deploy editssaves tolabelssaves to
@@
Figure 1. Before a developer deploys a prompt template to production, spade analyzes the deltas (i.e., diffs) between consecutive prompt templates to generate assertions. Then, spade uses labeled pipeline inputs and outputs to filter out redundant and inaccurate assertions, while maintaining coverage of bad outputs.
There is a lot of excitement around the use of large language models (LLMs) for processing, understanding, and generating data (Fernandez et al., 2023). Without needing large labeled datasets, one can easily create an LLM pipeline for any task on a collection of data items—simply by crafting a natural language prompt that instructs the LLM on what to do with each item. This could span traditional data processing tasks, such as summarizing each document in a corpus, extracting entities from each news article in a collection of articles, or even imputing missing data for each tuple in a relation (Narayan et al., 2022). LLMs additionally enable new and more complex data processing tasks that involve generating data, e.g., an LLM could write an explanation for why a product was recommended to a user, author emails to potential sales leads, or draft blog posts for marketing and outreach. In all of these data processing tasks, deploying these LLM pipelines at scale—either offline, on each batch of data items, or online, as and when new items arrive—presents significant challenges, due to data quality errors made by LLMs seemingly at random (Kalai and Vempala, 2023)—with LLMs often disregarding instructions, making mistakes with the output format, or hallucinating facts (Zamfirescu-Pereira et al., 2023; Shankar et al., 2022a).
One approach to catch errors in deployed LLM pipelines is via data quality assertions. Indeed, multiple recent papers (Rebedea et al., 2023; Singhvi et al., 2023; Kim et al., 2023) and LLM pipeline authoring systems (Guardrails, 2023; Langchain, 2023; Llama Index, 2023) provide mechanisms to embed manually-provided or selected assertions as part of LLM pipelines to catch errors during deployment. However, determining which assertions to add remains an open problem—and is a big customer painpoint based on our experience at LangChain—a company that helps people build LLM pipelines. Developers often find it difficult to determine the right set of assertions for their custom tasks (Parnin et al., 2023). Challenges include predicting all possible LLM failure modes, the time-consuming nature of writing assertions with various specification methods (like Python functions or LLM calls), the necessity for precision in assertions (especially those involving LLM calls), and the fact that many LLM pipeline developers lack software engineering expertise or coding experience (Kim et al., 2023; Zamfirescu-Pereira et al., 2023). Moreover, if there are non-informative assertions or too many of them, developers can get overwhelmed monitoring the results of these assertions. While there is some work on automatically detecting errors in traditional ML pipelines (Schelter et al., 2018; Shankar et al., 2023a; Polyzotis et al., 2018; Breck et al., 2019a), this line of work operates on many structured records at a time, and doesn’t apply to an unstructured setting. So, we target the following question: can we identify data quality assertions for LLM pipelines with as little effort from developers as possible?
@@
Example LLM Pipeline. Consider an LLM pipeline for a movie streaming platform, where the task is to generate a paragraph of text explaining why a specific movie was recommended to a specific user. A developer might write a prompt template like: “Given the following information about the user, {personal_info}, and information about a movie, {movie_info}: write a personalized note for why the user should watch this movie” to be executed for many user-movie pairs. In theory, this prompt seems adequate, but the developer might observe some data quality issues while testing it across different inputs: the LLM output might reference a movie the user never watched, cite a sensitive attribute (e.g., race or ethnicity), or even exhibit a basic issue by being too short. To fix these problems, developers typically add instructions to the prompt to catch these data quality issues, such as “Don’t reference race in your answer”. However, the LLM may still violate these instructions in an unpredictable manner for some data items, requiring assertions applied to LLM outputs post-hoc, during deployment.
Analyzing Prompt Version Histories. To automatically synthesize assertions for developers, our first insight is that we can mine prompt version histories to identify assertion criteria for LLM pipelines, since developers implicitly embed data quality requirements through changes to the prompt, or prompt deltas, over time. In our example above, instructions such as “Make sure your response is at least 3 sentences” or “Don’t reference race in your answer” could each correspond to a candidate assertion. To understand the types of prompt deltas in LLM pipelines and verify their usefulness for data quality assertions, we present an analysis of prompt version histories of 19 custom pipelines from LangChain users. Using this analysis, we construct a taxonomy of prompt deltas (Figure 2), which may be of independent interest for researchers and practitioners studying how to best build LLM pipelines.
Redundancy in Data Quality Assertions. Our second insight from analyzing these pipelines is that if we were to create candidate assertions from prompt deltas, say, using an LLM, there may be far too many assertions—often exceeding 50 for just a few prompt deltas. Many of these assertions (or equivalently, prompt deltas) are redundant, while some are too imprecise or ambiguous to be useful (e.g., “return a concise response”). Reducing this redundancy is nontrivial, even for engineers: assertions themselves can involve LLM calls with varying accuracies, motivating an automated component that can filter out bad candidates. One approach is to use a small handful of developer-labeled LLM outputs to estimate each data quality assertion’s false positive and false negative rate—but collectively determining the right set of assertions that can catch most errors, while incorrectly flagging as few outputs as possible, is not straightforward.
spade: Our Data Quality Assertion Generation Framework. In this paper, we leverage the aforementioned insights to develop spade (System for Prompt Analysis and Delta-Based Evaluation, Figure 1). spade’s goal is to select a set of boolean assertions with minimal overlap, while maximizing coverage of bad outputs and minimizing false failures (correct outputs that are incorrectly flagged) for the conjunction of selected assertions. We decompose spade into two components—candidate assertion generation and filtering.
Component 1: Prompt Deltas for Candidate Assertion Generation. For generating candidate assertions, instead of directly querying an LLM to “write assertions for x prompt,” which causes the LLM to miss certain portions of the prompt, we generate candidates from each prompt delta, which typically indicate specific failure modes of LLMs. spade leverages the aforementioned taxonomy of prompt deltas we constructed by first automatically categorizing deltas using the taxonomy, then synthesizing Python functions (that may include LLM calls) as candidate assertions. At LangChain, we publicly release this component of spade—which has been subsequently used for over 2000 pipelines across more than 10 sectors like finance, medicine, and IT (Shankar et al., 2023b). We present an analysis of this usage in Section 2.4.
Component 2: Filtering Candidate Assertions with Limited Data. To filter out incorrect and redundant candidate assertions, instead of requiring cumbersome manual selection or even fine-tuning of separate models (Saad-Falcon et al., 2023; Wang et al., 2023b), we propose an automated approach that only requires a small handful of labeled examples, which are usually already present in most target applications. Using these examples, we could estimate each assertion’s false failure rate (FFR), i.e., how often an assertion incorrectly flags failures, and eliminate individual assertions that exceed a given threshold. However, given we are selecting a set of assertions, the set may still exceed the FFR threshold and flag too many failures incorrectly, and redundancies may persist. We show that selecting a small subset of assertions to meet failure coverage and FFR criteria is NP-hard. That said, we may express the problem as an integer linear program (ILP) and use an ILP solver to identify a solution in a reasonable time given the size of our problem (hundreds to thousands of variables).
In some cases, when there are limited developer-provided examples, we find that labeled LLM outputs may not cover all failure modes, leading to omission of valuable data quality assertions. For instance, in our movie recommendation scenario, an assertion that correctly verifies if the output is under 200 words will get discarded if all outputs in our developer-labeled sample respect this limit. To expand coverage, active learning and weak supervision approaches can be used to sample and label new LLM input-output pairs for each candidate assertion (Castro and Nowak, 2008; Ratner et al., 2020), but this may be expensive or inaccessible for non-programmers. We introduce assertion subsumption as a way of ensuring comprehensive coverage: if one assertion doesn’t encompass the failure modes of another, both may be selected. As such, spade selects a minimal set of assertions that respects failure coverage, accuracy, and subsumption constraints.
Overall, we make the following contributions:
  • We identify prompt version history as a rich source for LLM output correctness, creating a taxonomy of assertion criteria from 19 diverse LLM pipelines (Section 2),
  • We introduce spade, our system that automatically generates data quality assertions for LLM pipelines. In a public release of spade’s candidate assertion generation component [^1] https://spade-beta.streamlit.app, we observe and analyze 2000+ deployments (Section 2.4),
  • We present a method to select a minimal set of assertions while meeting coverage and accuracy requirements, used by spade to reduce the number of assertions. For low-data settings, we introduce assertion subsumption as a novel proxy for coverage (Section 3), and
@@
  • We demonstrate spade’s effectiveness on nine real-world LLM pipelines (eight of which we open-source). In our low-data setting (approximately 75 inputs and outputs per pipeline), our subsumption-based solution outperforms simpler baselines that do not consider interactions between assertions by reducing the number of assertions by 14% and lowering the false failure rate by 21% (Section 4).
@@ 2. Identifying Candidate Assertions
Version $i$ $\Delta\mathcal{P}_{\_}i$ $\Delta\mathcal{P}_{\_}i$ Category Possible New Assertion Criteria
1 + Given the following information about the user, {personal_info}, and information about a movie, {movie_info}: write a personalized note for why the user should watch this movie. Inclusion Response should be personalized and relevant to the given user information
2 + Include elements from the movie’s genre, cast, and themes that align with the user’s interests. Inclusion Response includes specific references to the user’s interests related to the movie’s genre, cast, and themes
3 + Ensure the recommendation note is concise. Qualitative Assessment Response should be concise
4 - Ensure the recommendation note is concise. + Ensure the recommendation note is concise, not exceeding 100 words. Count Response should be within the 100 word limit
5 - Include elements from the movie’s genre, cast, and themes that align with the user’s interests. + Mention the movie’s genre and any shared cast members between the {movie_name} and other movies the user has watched. Inclusion Response should mention genre and verify cast members are accurate
6 + Mention any awards or critical acclaim received by movie_name. Inclusion Response should include references to awards or critical acclaim of the movie
7 + Do not mention anything related to the user’s race, ethnicity, or any other sensitive attributes. Exclusion Response should not include references to sensitive personal attributes
Table 1. Comparison of 7 prompt versions for an LLM pipeline to write personalized movie recommendations. In each $\Delta\mathcal{P}_{\_}i$, a sentence starts with “+” if it is a newly added sentence; a sentence starts with “-” if it is removed from $\mathcal{P}_{\_}{i-1}$. For each $\Delta\mathcal{P}_{\_}i$, we list its category (using the taxonomy in Figure 2) and possible new assertion criteria.
@@ 2.1. Prompt Deltas
A single-step LLM pipeline consists of a prompt template $\mathcal{P}$, which is formatted with a serialized input tuple $t$ to derive a prompt that is fed to an LLM, returning a response. There can be many versions of $\mathcal{P}$, depending on how a developer iterates on their prompt template. Let $\mathcal{P}_{\_}0$ be the empty string, the 0th version, and let $\mathcal{P}_{\_}i$ be the ith version of a template. In the movie recommendation example from the introduction, suppose there are 7 versions, where $\mathcal{P}_{\_}7$ is the following: “Given the following information about the user, {personal_info}, and information about a movie, {movie_info}: write a personalized note for why the user should watch this movie. Ensure the recommendation note is concise, not exceeding 100 words. Mention the movie’s genre and any shared cast members between the {movie_name} and other movies the user has watched. Mention any awards or critical acclaim received by {movie_name}. Do not mention anything related to the user’s race, ethnicity, or any other sensitive attributes.”
We define a prompt delta $\Delta\mathcal{P}_{\_}{i+1}$ to be the diff (or difference) between $\mathcal{P}_{\_}i$ and $\mathcal{P}_{\_}{i+1}$. Concretely, a prompt delta $\Delta\mathcal{P}$ is a set of sentences, where each sentence is tagged as an addition (i.e., “+”) or deletion (i.e., “-”). Table 1 shows the $\Delta\mathcal{P}$s for a number of versions for our example. Each sentence in $\Delta\mathcal{P}_{\_}i$ is composed of additions (i.e., new sentences in $\mathcal{P}_{\_}i$ that didn’t exist in $\mathcal{P}_{\_}{i-1}$) and deletions (i.e., sentences in $\mathcal{P}_{\_}{i-1}$ that don’t exist in $\mathcal{P}_{\_}{i}$). A modification to a sentence is represented by a deletion and addition—for example, $\Delta\mathcal{P}_{\_}{6}$ in Table 1 contains some new instructions added to a sentence from $\mathcal{P}_{\_}{5}$. Each addition in $\Delta\mathcal{P}_{\_}{i}$ indicates possible assertion criteria, as shown in the right-most column of Table 1.
@@ 2.2. Prompt Delta Analysis
To understand what assertions developers may care about, we turn to real-world LLM pipelines. We analyzed 19 LLM pipelines collected from LangChain users, each of which consists of between three and 11 historical prompt template versions. These pipelines span various tasks across more than five domains (e.g., finance, marketing, coding, education, health), from generating workout summaries to a chatbot acting as a statistics tutor. Table 7 in Appendix B shows a summary of the pipelines, including a description of each pipeline and the number of prompt versions. For each pipeline, we categorized prompt deltas, i.e., $\Delta\mathcal{P}_{\_}i$, into different types—for example, instructing the LLM to include a new phrase in each response (i.e., inclusion), or instructing the LLM to respond with a certain tone (i.e., qualitative criteria). Two authors iterated on the categories 4 times through a process of open and axial coding, ultimately producing the taxonomy in Figure 2. The taxonomy-annotated dataset of prompt versions can be found online [^2] https://github.com/shreyashankar/spade-experiments/blob/main/taxonomylabels.csv.
We divide deltas into two main high-level categories: Structural and Content-Based. Around 35% categories identified across all deltas in our dataset were Structural, and 65% were Content-Based. Structural deltas indicate a minor restructuring of the prompt, without changing any criteria of a good response (e.g., adding a newline for readability), or specification of the intended output (e.g., JSON or Markdown). Plausible assertion criteria based on structural deltas would check if the LLM output adheres to the user-specified structure. On the other hand, content-based deltas indicate a change in the meaning or definition of the task. Content-based deltas include descriptions of the workflow steps that the LLM should perform (e.g., “first, do X, then, come up with Y”), instructions of specific phrases to include or exclude in responses, or qualitative indicators of good responses (e.g., “maintain a professional tone”). The Data Integration subcategory (under Content-Based deltas) concerns adding new sources of context to the prompt—for example, adding a new variable like “{movie_info}” to the prompt, indicating a new type of information to be analyzed along with other content in the prompt. For some illustrative examples of prompt deltas for each category, we categorize the prompt deltas in Table 1, and in Table 2, we show sample prompt deltas for each category in our taxonomy. This taxonomy may be of independent interest to researchers studying the process of prompt engineering, as well as practitioners seeking to identify ways to improve their prompts for production LLM pipelines.
Overall, our exercise in building this taxonomy reveals two key findings. First, developers across diverse LLM pipelines iterate on prompts in similar ways as encoded as nodes in the taxonomy, many of which correspond to aspects that can be explicitly checked via assertions. Second, we find that an automated approach to synthesize data quality assertions may be promising, since our taxonomy reveals several such instances where the prompt delta could directly correspond to an assertion that captures the same requirement. For example, many deltas correspond to instructions to include or exclude specific phrases, indicating that developers may benefit from assertions that explicitly verify the inclusion or exclusion of such phrases in LLM outputs.
However, if we are to use an LLM to automatically generate assertion criteria based on our taxonomy of deltas, we also need to test whether LLMs can accurately identify the categories corresponding to a delta. Therefore, we confirmed GPT-4’s correct categorization of prompt deltas (as of October 2023): we assigned ground truth categories to all prompt deltas from the 19 pipelines, and GPT-4 achieved an F1 score of 0.8135. The prompt used for category extraction from prompt deltas is detailed in Appendix C.
@@
Response Format Instruction Structure guidelines. “+ Start response with ‘You might like…’.”
Example Demonstration Illustrative example. “+ For example, here is a response for sci-fi fans…”
Prompt Clarification Refines prompt/removes ambiguity. “- Discuss/+ Explain movie fit…”
Workflow Description Describe “thinking” process. “+ First, analyze viewing history…”
Data Integration Adds placeholders. “+ Include user’s {genre} reviews.”
Quantity Instruction Adds numerical content. “+ Keep note under 100 words.”
Inclusion Instruction Directs specific content. “+ Mention movie awards/acclaim.”
Exclusion Instruction Advises on omissions. “+ Avoid movie plot spoilers.”
Qualitative Criteria Sets stylistic attributes. “+ Maintain friendly, positive tone.”